OpenStudy (anonymous):

How do i solve a problem like: Ryan can dig a hole in 5 hours, it takes Sue 6 hours to dig the same hole. How long will it take both of them together?

OpenStudy (anonymous):

\[\frac{x}{5}+\frac{x}{6}=1\]

OpenStudy (anonymous):

\[5x+6x=30\]

OpenStudy (anonymous):

\[11x=30\]\[x=\frac{30}{11}=2.72\]

OpenStudy (anonymous):

2.73 hours

OpenStudy (anonymous):

ty a lot

OpenStudy (anonymous):

Factor completely 5y\[^{3}\] - 125 y

OpenStudy (anonymous):

5y (raised to the third) - 125y